What is the measure of an angle whose supplement is 3 times its complement
Answer by mananth(16946)

measure of angle = x
supplement = 180-x
complement = 90-x
..
180-x = 3(90-x)
180-x = 270-3x
3x-x=270-180
2x=90
/2
x= 45 the measure of the angle
..
m.ananth@hotmail.ca